# Mixed-Integer Programming STALC Stratified Topological Autonomy for Long-Range Coordination

### v1

@@ VII-A Robot Route Assignment
0:
$M$ List of routes for each robot ($M_{a}~{}\forall a$)
1:  $M\leftarrow$ NewList()
2:  for all $a\in[1,n_{A}]$ do
3:     $M_{a}\leftarrow$ NewList()
4:     $s_{a}\leftarrow$ RemoveItem($S$)
5:     $M_{a}\leftarrow$ AddItem($s_{a}$)
6:     $p_{s_{a},1}\leftarrow p_{s_{a},1}-1$
7:     $(n_{1},n_{2})\leftarrow s_{a}$
8:     for all $t\in[2,n_{T}]$ do
9:        for all $l\in L$ do
10:           if $p_{l,t}>0$ then
11:              $(v_{1},v_{2})\leftarrow l$
12:              if $n_{2}=v_{1}$ then
13:                 $M_{a}\leftarrow$ AddItem($l$)
14:                 $p_{l,t}\leftarrow p_{l,t}-1$
15:                 $(n_{1},n_{2})\leftarrow(v_{1},v_{2})$
16:                 break
17:              end if
18:           end if
19:        end for
20:     end for
21:     $M\leftarrow$ AddItem($M_{a}$)
22:  end for
23:  return  $M$
@@ VII-B Mid-Range Plan Generation
When a team of robots is commanded to navigate an edge in the graph, we arbitrarily designate a robot to be the leader of the team and all other robots are considered followers. Each robot is then assigned a position to maintain in the formation for that edge relative to the team leader. Robots within a team share their positions and planned low-level trajectories with team members to enable predictive collision avoidance.
Based on their role in the team, we provide each robot with a mid-range plan to follow that minimizes visibility to potential observers, which we computed with A in Section V-B8. For each robot, we continuously truncate the path to start at the closest point to the robot’s current position. Additionally, we truncate the end of the leader’s path to a receding horizon goal based on its maximum speed and planning time horizon. This receding horizon goal is shared with the follower robots. We then truncate each follower’s path to its own receding horizon goal, which is calculated based on the leader’s receding horizon goal and the follower’s assigned position in the formation.
@@ VIII Low-Level Planner
After we generate a mid-level plan, it is then executed by the low-level planner, which runs decentralized on each robot.
To avoid obstacles, which may not exist in the global map, we leverage TerrainNet [40] for local mapping. TerrainNet takes RGB-D images of the environment as input and produces semantic and elevation maps as output. We produce a costmap, $\mathcal{C}$, by assigning a cost, $[0,255]$, to each semantic label, which correspond to the features of the environment (e.g., bush, rock, tree, dirt, etc.) that occupy the terrain.
@@
We represent the robots using Dubins Car kinematics with bounded wheel velocity and acceleration as follows.
$\begin{split}\mathbf{x_{t}}=&\ \begin{bmatrix}\mathbf{p_{t}}\ \theta_{t}\end{bmatrix}^{T}\\ \mathbf{u_{t}}=&\ \begin{bmatrix}v_{t}\ \alpha_{t}\end{bmatrix}^{T}\\ \mathbf{w_{t}}=&\begin{bmatrix}\frac{1}{r}&\frac{b}{2r}\\ \frac{1}{r}&-\frac{b}{2r}\end{bmatrix}\mathbf{u_{t}}\\ \mathbf{w^{max}_{t}}=&\ \max(\mathbf{w_{t}},\frac{v_{min}}{r},\mathbf{w_{t-1}}-\frac{a_{max}}{r}\Delta t)\\ \mathbf{w^{clamp}_{t}}=&\ \min(\mathbf{w^{max}_{t}},\frac{v_{max}}{r},\mathbf{w_{t-1}}+\frac{a_{max}}{r}\Delta t)\\ \mathbf{u^{clamp}_{t}}=&\ \begin{bmatrix}\frac{r}{2}&\frac{r}{2}\\ \frac{r}{b}&-\frac{r}{b}\end{bmatrix}\mathbf{w^{clamp}_{t}}\\ \mathbf{x_{t+1}}=&\ \mathbf{x_{t}}+\begin{bmatrix}\cos(\theta_{t})&0\\ \sin(\theta_{t})&0\\ 0&1\end{bmatrix}\mathbf{u^{clamp}_{t}}\Delta t\end{split}$ (33)
@@
Here $\mathbf{x_{t}}$ is the state, $\mathbf{p_{t}}$ is the position, $\theta_{t}$ is the orientation, $\mathbf{u_{t}}$ is the control, $v_{t}$ is the linear velocity, $\alpha_{t}$ is the angular velocity, $r$ is the wheel radius, $b$ is the wheel base, $\mathbf{w}_{t}$ is the wheel velocities, $v_{min}$ is the minimum linear velocity, $v_{max}$ is the maximum linear velocity, $a_{max}$ is the maximum linear acceleration, and $\Delta t$ is the discrete time step.
@@
For local planning, we use MPPI [33]. We design our cost function to be a weighted sum of cost components to encourage: navigating towards the goal in the graph (through terms for goal distance $q_{d}$, goal heading $q_{h}$, and goal pointing $q_{p}$), following the mid-range path (through terms for mid-range path distance $q_{md}$ and mid-range path heading $q_{mh}$), and avoiding obstacle collisions, inter-robot collisions, and terrain that cannot be traversed (through terms for a costmap $q_{c}$ and trajectory collision $q_{tc}$). Our overall cost function is as follows.
$q(\mathbf{x_{t}})=$ $\ \gamma_{d}q_{d}(\mathbf{x_{t}})+\gamma_{h}q_{h}(\mathbf{x_{t}})+\gamma_{p}q_{p}(\mathbf{x_{t}})+\gamma_{md}q_{md}(\mathbf{x_{t}})$
$+\gamma_{mh}q_{mh}(\mathbf{x_{t}})+q_{c}(\mathbf{x_{t}})+q_{tc}(\mathbf{x_{t}})$ (34)
@@
In this cost function, $\gamma_{d}$, $\gamma_{h}$, $\gamma_{p}$, $\gamma_{md}$, and $\gamma_{mh}$ are heuristically selected cost weights.
@@
We define the difference between the current state of the robot $\mathbf{x_{t}}$ and the goal state $\mathbf{x^{g}}$, the mid-range path $\mathbf{x^{m}_{t}}$, and the previous state $\mathbf{x_{t-1}}$. Additionally, we define $\Delta\mathbf{x^{i}_{t^{i}t}}$, the difference between the current state of the robot and every state in the planned trajectory of each other robot on the team, $\mathbf{x^{i}_{t^{i}}}$. Here, $i$ is the robot index and $t^{i}$ is the time index along robot $i$’s trajectory.
$\mathbf{\Delta x^{g}_{t}}=$ $\ \mathbf{x^{g}}-\mathbf{x_{t}}$ (35)
$\mathbf{\Delta x^{m}_{t}}=$ $\ \mathbf{x^{m}_{t}}-\mathbf{x_{t}}$ (36)
$\mathbf{\Delta x_{t}}=$ $\ \mathbf{x_{t}}-\mathbf{x_{t-1}}$ (37)
$\mathbf{\Delta x^{i}_{t^{i}t}}=$ $\ \mathbf{x^{i}_{t^{i}}}-\mathbf{x_{t}}$ (38)
@@
Finally, we can define the cost components. We use parameters for goal pointing radius $r_{p}$, the maximum mid-range path cost distance $d_{m}$, the lethal costmap penalty $c$, the lethal costmap threshold $\ell$, the trajectory collision radius $r_{t}$, and the planning time horizon $T$.
$q_{d}(\mathbf{x_{t}})=$ $\ \|\mathbf{\Delta p^{g}_{t}}\|$ (39)
@@
$q_{h}(\mathbf{x_{t}})=$ $\ \exp\left(\left(\frac{\mathbf{\Delta p^{g}_{t}}}{2}\cdot\begin{bmatrix}\cos(\theta^{g})\\ \sin(\theta^{g})\end{bmatrix}\right)^{2}\right)$
$\cdot\min(-\cos(\Delta\theta^{g}_{t}),0.9)$ (40)
$q_{p}(\mathbf{x_{t}})=$ $\ \mathbb{I}_{\{\|\mathbf{\Delta p^{g}_{t}}\|<r_{p}\}}$
$\cdot\max(-\cos(\operatorname{atan2}(\mathbf{\Delta p^{g}_{t}})-\theta^{g}),0)$ (41)
$q_{md}(\mathbf{x_{t}})=$ $\ \min\left(\frac{\|\mathbf{\Delta p^{m}_{t}}\|}{d_{m}},1.0\right)^{2}$ (42)
$q_{mh}(\mathbf{x_{t}})=$ $\ \left(1-\frac{\mathbf{\Delta p_{t}}}{\|\mathbf{\Delta p_{t}}\|}\cdot\begin{bmatrix}\cos(\theta^{m})\\ \sin(\theta^{m})\end{bmatrix}\right)$ (43)
$q_{c}(\mathbf{x_{t}})=$$\ \begin{cases}1\mathrm{e}{10},&\text{if }\mathcal{C}(\mathbf{p_{t}})\geq\ell\text{ and }\mathcal{C}(\mathbf{p_{t-1}})<\ell\\ 1\mathrm{e}{10},&\text{else if }\mathcal{C}(\mathbf{p_{t}})\geq\ell\text{ and }t=T\\ c,&\text{else if }\mathcal{C}(\mathbf{p_{t}})\geq\ell\text{ and }\mathcal{C}(\mathbf{p_{t-1}})\geq\ell\\ \frac{\mathcal{C}(\mathbf{p_{t}})}{\ell}&\text{otherwise }\end{cases}$ (44)
$q_{tc}(\mathbf{x_{t}})=$ $\ \mathbb{I}_{\left\{\bigvee_{i,t^{i}}\left(\|\mathbf{\Delta p^{i}_{t^{i}t}}\|<r_{t}\right)\right\}}\cdot 1\mathrm{e}{10}$ (45)
For our experiments with Clearpath Warthogs, we used the following parameters: $r=0.279m$, $b=1.175m$, $v_{min}=-0.5\frac{m}{s}$, $v_{max}=2.0\frac{m}{s}$, $a_{max}=4.0\frac{m}{s^{2}}$, $r_{p}=1m$, $d_{m}=10m$, $c=30$, $\ell=200$, $\gamma_{d}=1$, $\gamma_{h}=0.1$, $\gamma_{p}=0.1$, $\gamma_{md}=5.0$, and $\gamma_{mh}=1.0$. For MPPI, we used $\Delta t=\frac{1}{15}s$, $K=3000$ samples, time horizon $T=5s$, temperature $\lambda=0.01$ and sampling variance of $\Sigma=1.94$.
@@ IX Results and Discussion
We first demonstrate our graph planning approach in simulated scenarios to illustrate the advantages of our approach. We then demonstrate the full hierarchical planning architecture in hardware experiments.
@@ IX-A1 Illustrative Example
Figure 9: MIP problem solution to an illustrative multi-robot reconnaissance scenario, sketched in (a). Ten robots start at node 1 with a goal of at least one robot reaching node 5 to observe the adversary units. The light green regions represent areas of cover. The color of the edges indicate the threat level for transitioning between nodes from low to high (LTH): blue, yellow, red. Edge (3,5) has the highest weight due to its visibility by the adversary. In each subplot, the edge labels indicate the edge cost under the current conditions and show in brackets the desired number of robots for traversal of the edge, $a_{e}$, due to vulnerability. Overwatch opportunities are shown with a pink arrow from the overwatch node pointing to the edge that can be monitored. At each time step, the position of each robot is shown.
@@
We consider the example in Fig. 9 which depicts a simple reconnaissance scenario. There are 10 robots starting at node 1 and the goal is for at least one robot to reach node 5 while minimizing visibility. Fig. 9 steps through a solution to this problem, which we generated using our MIP optimization problem. In Fig. 9(a), the edge costs are shown as the maximum cost for one robot to traverse without any overwatch, teaming, or vulnerability cost reductions. The edge labels in each subplot show the edge cost and the number of robots desired in brackets ([]). The color of the edge indicates threat level blue (lowest), yellow, or red (highest). To encourage teaming, each additional robot on an edge reduces the cost by 1. Both directions of edges (1,3), (3,5), and (4,5) are considered vulnerable and at least four robots are desired; the cost reduction for each robot up to four is 10. Overwatch opportunities are indicated by the pink arrows. Overwatch can be provided from node 2 for both directions of edge (2,4) and node 3 for both directions of edge (4,5), which can reduce the cost by up to 20 or 60, respectively, when two robots are providing overwatch. Each additional robot providing overwatch after the first two would reduce the cost by 2. The dynamic costs incurred by each edge are updated in the graphs in Fig. 9 to reflect the teaming, vulnerability, and overwatch constructs being utilized. In this example, we scale the time cost by 10 to encourage reaching the goal in minimal time. The total accumulated cost is tracked in the captions.
In this solution, we see the robots break into three teams to work together for a subset of the robots to safely traverse to the target node. The robots all move together on edge (1,2) to get the largest benefit from teaming. Two robots then remain at node 2 to provide overwatch to the four robots traversing edge (2,4) and the remaining four continue toward node 3 to be prepared to provide overwatch in the next step. The greatest benefits from overwatch at nodes 2 and 3 are with 2 robots. Four robots (rather than two) go to node 3 to also receive a larger teaming benefit on edge (2,3) and additional rewards for overwatch in the next step. It would have been equivalent cost for two of those robots to continue on with robots 4-7. In time step 3, robots 4-7 traverse a vulnerable edge, getting the largest reward for having four robots and enabling moving in a protective formation. In the last step, robots 4-7 reach the goal node. The cost does not increase since the robots moved to a node in this step and there is no longer a time cost since there are no robots remaining on edges.
By solving our problem to optimality, we are guaranteed to have minimized the total cost. However, this depends greatly on the weights/cost assigned to the problem and the prioritization of minimizing traversing cost versus minimizing time. For example, if we did not scale our time cost by 10, the resulting solution would keep more robots together to receive more teaming benefits for moving together and from additional overwatch cost reductions. As in Fig. 9(b), all robots would go to node 2. Then the robots would split into only two teams: a team for overwatch that moves from node 2 to 3 and the team being monitored on edges (2,4) and (4,5). Ultimately, requiring one additional time step.

### v2

@@ VII-A Robot Route Allocator
2:
$M$ List of routes for each robot ($M^{i}~\forall i\in[1,n_{A}]$)
3:for all $i\in[1,n_{A}]$ do
[…]
6:for all $t\in[2,n_{T}]$ do
7:  for all $l\in L$ do
8:   if $p_{l,t}>0$ and $l\in A(s^{i}_{t-1})$ then
9:    $M^{i}\leftarrow$ AddItem($l$);  $p_{l,t}\leftarrow p_{l,t}-1$;  $s^{i}_{t}\leftarrow l$
10:    break
11:return $M$
@@ VII-B Mid-Range Plan Generation
When a coalition of robots, $\mathcal{C}_{e}$ is assigned to traverse an edge in the graph, members are ordered alphabetically; the first robot serves as leader, while the others are sequentially assigned follower positions. In this paper, we employ a leader-follower line formation, where each robot follows the one preceding it. Each robot’s route, from Algorithm 3, maps to an edge-specific path $\mathcal{T}_{v_{i}v_{j}}$ that minimizes visibility to potential observers, which we computed with A in Sec. VI-B7. The mid-range plan for a robot, $\mathcal{T}^{i}$, starts at the closest point along $\mathcal{T}_{v_{i}v_{j}}$ to the robot’s current position and truncates at a receding horizon goal along $\mathcal{T}_{v_{i}v_{j}}$. For the leader, this goal is based on its maximum speed and planning time horizon. For the follower, this goal is based on a set following distance from the planned goal of the robot preceding it in the formation. Thus, the coalition’s set of mid-range plans $\mathcal{T}^{\mathcal{C}_{e}}$ is a function of $\mathcal{T}_{v_{i}v_{j}}$ and $X_{B}^{\mathcal{C}_{e}}$.
@@ VIII Low-Level Planner
As described in Sec. III, we solve a local optimization problem of the form in (1) to execute edge-specific multi-robot motion planning. Rather than solving the optimization problem over the joint state space, we leverage a distributed receding-horizon NMPC control approach [63], where coupled costs and constraints are approximated by robot team members sharing state and policy information from the prior planning interval. Given the set of mid-range plans $\mathcal{T}^{{C}_{e}}$ generated in VII-B, we can also use a short, computationally tractable time-horizon and a receding-horizon goal $X^{\mathcal{C}_{e}}_{D}$ constructed from the set of terminal points in $\mathcal{T}^{{C}_{e}}$. Our approach uses only a state-cost $g_{x,e}$, and does not include an input cost $g_{u,e}$, or state or input constraints $c_{x_{e}}$ or $c_{u_{e}}$.
@@
For the dynamics $f(x_{t})$ in (1), we use Dubins Car kinematics with bounded wheel velocity and acceleration:
$x_{t}$ $=\begin{bmatrix}p_{t}\ \theta_{t}\end{bmatrix}^{T},\kern 5.0ptu_{t}=\begin{bmatrix}v_{t}\ \alpha_{t}\end{bmatrix}^{T},\kern 5.0ptw_{t}=\begin{bmatrix}\frac{1}{r}&\frac{b}{2r}\\ \frac{1}{r}&-\frac{b}{2r}\end{bmatrix}u_{t}$
$w^{\textrm{max}}_{t}$ $=\max(w_{t},\frac{v_{\textrm{min}}}{r},w_{t-1}-\frac{a_{\textrm{max}}}{r}\Delta t)$
$w^{\textrm{clamp}}_{t}$ $=\min(w^{\textrm{max}}_{t},\frac{v_{\textrm{max}}}{r},w_{t-1}+\frac{a_{\textrm{max}}}{r}\Delta t)$ (35)
$x_{t+1}$ $=x_{t}+\begin{bmatrix}\cos(\theta_{t})&0\\ \sin(\theta_{t})&0\\ 0&1\end{bmatrix}\begin{bmatrix}\frac{r}{2}&\frac{r}{2}\\ \frac{r}{b}&-\frac{r}{b}\end{bmatrix}w^{\textrm{clamp}}_{t}\Delta t.$
@@
Here we define the state $x_{t}\in\mathbb{R}^{3}$, position $p_{t}\in\mathbb{R}^{2}$, orientation $\theta_{t}$, control $u_{t}\in\mathbb{R}^{2}$, linear velocity $v_{t}$, angular velocity $\alpha_{t}$, wheel radius $r$, wheel base $b$, wheel velocities $w_{t}\in\mathbb{R}^{2}$, minimum linear velocity $v_{\textrm{min}}$, maximum linear velocity $v_{\textrm{max}}$, maximum linear acceleration $a_{\textrm{max}}$, and time step $\Delta t$.
@@
Since the cost functions from (1) have been decoupled, for a given robot $i\in\mathcal{C}_{e}$, we can now write $g_{x}(X^{\mathcal{C}_{e}}_{t})$ as $g_{x,e}(x^{i}_{t})$. Dropping the $i$ for clarity, we design our cost function $g_{x,e}(x_{t})$ to be a weighted sum of cost components to encourage navigating towards the goal (distance $g_{\delta}$, heading $g_{h}$, and pointing $g_{p}$), following the mid-range path (distance $g_{m\delta}$ and heading $g_{mh}$), avoiding untraversable terrain (terrain costmap $g_{c}$), and inter-robot collisions (trajectory collision $g_{tc}$).
To avoid obstacles, which may not exist in the global map, we leverage TerrainNet [64] for local mapping. TerrainNet takes RGB-D images of the environment as input and produces semantic and elevation maps as output. We produce a local costmap, $C_{l}$, by assigning a cost, $[0,255]$, to each semantic label, which correspond to the features of the environment (e.g., bush, rock, tree, dirt, etc.) that occupy the terrain. This approach aids in traversing diverse terrain, but could be substituted for another local mapping approach. The overall cost function is as follows:
$g_{x,e}(x_{t})=$ $\ \gamma_{\delta}g_{\delta}(x_{t})+\gamma_{h}g_{h}(x_{t})+\gamma_{p}g_{p}(x_{t})+\gamma_{m\delta}g_{m\delta}(x_{t})$
$+\gamma_{mh}g_{mh}(x_{t})+\gamma_{c}g_{c}(x_{t})+\gamma_{tc}g_{tc}(x_{t}).$ (36)
@@
In this cost function, $\gamma_{\delta}$, $\gamma_{h}$, $\gamma_{p}$, $\gamma_{m\delta}$, $\gamma_{mh}$, $\gamma_{c}$, and $\gamma_{tc}$ are heuristically selected cost weights.
@@
We define the differences between the current state of the robot $x_{t}$ and the following: the goal state $x_{D}$, the mid-range path $x^{m}_{t}$, and the previous state $x_{t-1}$. Additionally, we define $\Delta x^{i}_{t^{i}t}$, the difference between the current state of the robot and every state in the planned trajectory of each other robot in the same coalition, $x^{i}_{t^{i}}$. Here, $i\in\mathcal{C}^{e}$ is the robot index and $t^{i}$ is the time index along robot $i$’s trajectory.
$\begin{split}\Delta x^{d}_{t}=&\ x_{D}-x_{t},\quad\Delta x^{m}_{t}=x^{m}_{t}-x_{t}\\ \Delta x_{t}=&\ x_{t}-x_{t-1},\quad\Delta x^{i}_{t^{i}t}=x^{i}_{t^{i}}-x_{t}.\end{split}$ (37)
@@
Finally, we can define the cost components. We use parameters for goal pointing radius $r_{p}$, the maximum mid-range path cost distance $\delta_{m}$, the lethal costmap penalty $c$, the lethal costmap threshold $\ell$, the trajectory collision radius $r_{t}$, and the planning time horizon $T$.
$g_{\delta}(x_{t})=$ $\ \|\Delta p^{d}_{t}\|$
@@
$g_{h}(x_{t})=$ $\ \exp\left(\left(\frac{\Delta p^{d}_{t}}{2}\cdot\begin{bmatrix}\cos(\theta^{d})\\ \sin(\theta^{d})\end{bmatrix}\right)^{2}\right)$
$\cdot\min(-\cos(\Delta\theta^{d}_{t}),0.9)$
$g_{p}(x_{t})=$ $\ \mathbb{I}_{\{\|\Delta p^{d}_{t}\|<r_{p}\}}$
$\cdot\max(-\cos(\operatorname{atan2}(\Delta p^{d}_{t})-\theta^{d}),0)$
$g_{m\delta}(x_{t})=$ $\ \min\left(\frac{\|\Delta p^{m}_{t}\|}{\delta_{m}},1.0\right)^{2}$ (38)
[…]
$g_{tc}(x_{t})=$ $\ \mathbb{I}_{\left\{\bigvee_{i,t^{i}}\left(\|\Delta p^{i}_{t^{i}t}\|<r_{t}\right)\right\}}\cdot 1\mathrm{e}{10}.$
To optimize the objective function subject to the dynamics constraints, we use MPPI [52], though this approach could be substituted for other methods to planning dynamically feasible paths on costmaps.
@@ IX Results and Discussion
We first evaluate our graph planning approach in simulated scenarios. We then demonstrate the full hierarchical planning architecture, STALC, in autonomous hardware experiments.
@@ IX-A1 Illustrative Example
Figure 8: MIP problem solution to an illustrative MCRP, sketched in (a). Ten robots start at node 1 with a goal of at least one robot reaching node 5 to observe the adversary units. The light green regions represent areas of cover. The color of the edges indicate the threat level for transitioning between nodes from low to high (LTH): blue, yellow, red. Edge (3,5) has the highest weight due to its visibility by the adversary. In each subplot, the edge labels indicate the edge cost under the current conditions and show in brackets the desired number of robots for traversal of the edge, $a_{e}$, due to vulnerability. Overwatch opportunities, which are used in time steps 2 and 3, are shown with a pink arrow from the overwatch node pointing to the edge that can be monitored. At each time step, the position of each robot is shown.
@@
Fig. 8 depicts a simple version of the MCRP in which 10 robots start at node 1 and at least one robot must reach node 5 while minimizing visibility. We generate a solution to this problem using our graph planner, as shown in Fig. 8. In Fig. 8(a), the edge costs are shown as the maximum cost for one robot to traverse without reductions from overwatch, teaming, or vulnerability. The number of robots desired for each edge is shown in brackets ([]). The color of the edge indicates threat level blue (lowest), yellow, or red (highest). To encourage teaming, each additional robot on an edge reduces the cost by 1. Both directions of edges (1,3), (3,5), and (4,5) are considered vulnerable and at least four robots are desired; the cost reduction for each robot up to four is 10. Overwatch opportunities are indicated by the pink arrows. Overwatch can be provided from node 2 for both directions of edge (2,4) and node 3 for both directions of edge (4,5), which can reduce the cost by up to 20 or 60, respectively, when two robots are providing overwatch. Each additional robot providing overwatch after the first two would reduce the cost by 2. The dynamic costs incurred on each edge, which are based on the state of the robot team, are updated in the graphs in Fig. 8. In this example, we scale the time cost by 10 to encourage reaching the goal in minimal time. The total accumulated cost is tracked in the captions. In this solution, we see the robots break into three teams to work together for a subset of the robots to safely traverse to the target node. By solving our problem to optimality, we are guaranteed to have minimized the total cost. However, this depends greatly on the weights/cost assigned to the problem and the prioritization of minimizing traversing cost versus minimizing time.